OpenStudy (anonymous):

Simplify the expression below completely ( x^2-2 x-3)/( x^2-3 x-4) Hint: Factor the numerator and the denominator and then cancel any common factors.

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

\[\huge \frac{( x^2-2 x-3)}{( x^2-3 x-4)} =\frac{( x^2-3 x+x-3)}{( x^2-4 x+x-4)}\] \[\huge =\frac{x( x-3)+1(x-3)}{x( x-4)+ 1(x-4)}\] \[\huge =\frac{( x-3)(x+1)}{( x-4)(x+ 1)}\] Here we can see that (x+1 ) is the common factor inn Nr. and Dr. so we can cancell it out. \[\huge =\frac{( x-3)}{( x-4)}\] is the required simplest form of the given expression. @cancherolugano